Circular No. 8481

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004ht–2004io Further to IAUC 8362, J. Adelman-McCarthy, F. DeJongh, J. Frieman,

H. Lampeitl, and D. Tucker, Fermilab; B. Dilday and R. Kessler, University of Chicago; J. Holtzman, New Mexico State University (NMSU); and M. Sako, Stanford University, on behalf of the SDSS Collaboration, report the discovery of 22 supernovae on g′- and r′-band images taken at the SDSS 2.5m telescope. Spectroscopy was obtained during 2004 Oct. 14–Dec. 11 with the ARC 3.5-m telescope (+ DIS) at Apache Point and the Hobby-Eberly Telescope (+ LRS) at McDonald Observatory. Imaging was obtained with the ARC 3.5-m and the NMSU 1-m telescopes. Date of maximum observed g′ flux and approximate g′ magnitude on that date are tabulated below, together with supernova position, type, and redshift. Discovery dates: Sept. 22, 2004ht–hy, 2004ia; 24, 2004hz; Oct. 8, 2004ib, 2004id, 2004if; 10, 2004ic, 2004ie; 16, 2004ig, 2004ih; Nov. 3, 2004ii, 2004il–io; 5, 2004ij, 2004ik.

SN  2004 UT  α_{2000}  δ_{2000}  g′ Type z
2004ht Sept. 22 1^{h}04^{m}48.^{s}67 − 0^{o}53′45′′.5 21.1 II 0.067
2004hu Sept. 22 1 15 02.42 + 0 15 42.9 19.3 Ia 0.049
2004hv Sept. 22 23 48 25.95 + 1 03 50.3 20.2 II 0.061
2004hw Sept. 22 23 51 17.35 + 1 13 39.4 20.1 Ia 0.060
2004hx Sept. 24 2 41 12.60 − 0 52 46.0 17.2 II 0.014
2004hy Sept. 24 21 06 04.25 + 1 12 57.0 20.3 II 0.058
2004hz Oct. 8 2 05 03.00 + 0 50 11.9 20.1 Ia 0.143
2004ia Oct. 8 2 18 02.30 − 0 33 32.4 19.5 Ia 0.143
2004ib Oct. 10 2 40 56.40 − 0 10 48.3 20.3 Ic 0.056
2004ic Oct. 10 20 40 48.05 + 1 02 34.2 20.5 II 0.093
2004id Oct. 16 0 06 29.64 − 0 28 22.3 19.6 Ia 0.144
2004ie Oct. 16 22 01 46.58 + 1 14 12.0 17.5 Ia 0.052
2004if Oct. 16 22 39 08.54 − 0 51 54.6 21.5 Ia 0.322
2004ig Oct. 24 0 05 51.79 − 0 59 45.1 20.6 Ia 0.183
2004ih Oct. 24 2 15 54.31 + 1 04 09.9 20.0 Ia 0.154
2004ii Nov. 3 0 16 51.80 − 0 24 56.9 20.7 Ia 0.197
2004ij Nov. 5 1 21 11.54 − 0 55 09.2 21.3 Ia 0.218
2004ik Nov. 5 1 44 55.51 − 0 05 58.1 21.3 Ia 0.169
2004il Nov. 5 3 07 02.69 − 0 00 40.1 19.5 Ia 0.108
2004im Nov. 5 23 43 07.78 − 0 07 55.5 19.8 Ia 0.14
2004in Nov. 11 1 30 26.06 + 0 41 32.3 20.3 Ia 0.161
2004io Nov. 20 21 23 37.39 − 0 28 00.6 19.7 Ia 0.150

2005 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green